### v1

@@ Abstract
Eastern Boundary Upwelling Systems (EBUS) are characterized by wind-triggered upwelling of deep waters along the coast. They are hotspots of biological productivity and diversity and therefore have a high economic, ecological and social importance. In the past, different methods using surface data have been used to estimate upwelling. Recently, the IPCC has suggested directly assessing vertical velocities as a promising method. We use this method to study the two Atlantic EBUS from CMIP6 models from the HadGEM3-GC3.1 and the CNRM6-CM6 family, for both the historical period and a high-emission future scenario with spatial resolutions in the ocean component ranging from 1°to 1/12°. The two major upwelling regions are divided in subregions depending on their seasonality. The vertical transport index shows similar values to a wind-derived Ekman index. Directly evaluating upwelling from transport processes further provides information about the depth of the upwelling, which has previously been identified as an important factor for nutrient availability. We show that depending on the subregion of the upwelling system different cell structures can be seen in terms of depth and distance to the coast of maximum velocities. When looking at possible future changes high interannual variability limits the significance of the trends but could indicate a poleward shift of the upwelling regions. A detailed comparison of the spatial structures and the distinction in subregions is important to explain contradictory trends in previous works.
@@ 1 Introduction
Within the context of climate change, considering the future of EBUS therefore becomes of primary importance. Given that the major drivers of the upwelling in EBUS are the along shore trade winds, the community has focused its efforts in considering projected wind changes. Two main hypotheses are usually considered: i) the Bakun hypothesis (Bakun, 1990) states that along-shore wind stress will increase in future climates because of the intensification of the inland low pressure systems on the eastern borders of these regions. Seawater heat capacity being higher than soil heat capacity, one expects the surface inland temperature to rise faster than the seawater one, and therefore the atmospheric thermal low over land to strengthen. ii) The second hypothesis states that there will be a poleward migration of the trade winds as the atmospheric Hadley cell expands poleward, driving in turn the poleward shift of the subtropical high pressure atmospheric cells found in the four ocean basins (Rykaczewski et al., 2015). While one expects a general increase of the volume of water that is being upwelled to the surface in the case of an overall increase of the trade winds (hypothesis i), along-shore regional disparities are expected to happen under the poleward migration hypothesis, with an intensification/decrease of the upwelling in the poleward/equatorward parts of the EBUS.
Gulev et al. (2021)’s recent analysis, based on historical datasets of the wind fields, concluded to the rejection of the Bakun hypothesis. Although the increase of surface temperature over the last century was higher over land (1.59 [1.34 to 1.83]°C) than over the ocean (0.88 [0.68 to 1.01]°C), this increase did not significantly impact the intensity of the upwelling favourable surface wind stress ($\tau_{upw}$). On the other hand, the poleward migration of the Hadley cell was confirmed, and found to be more pronounced in the Northern than in the Southern Hemisphere. Additionally, contradicting trends were found for different EBUS. However, as already shown in the work of Sydeman et al. (2014) and Varela et al. (2015), trends appear to be very sensitive to the data product, period and geographical boundaries of the region of interest. This sensitivity highlights the lack of long-term robust observational datasets that can be used by the community to infer trends in upwelling systems.
Further, the ocean circulation, forced in part by the wind stress curl, is known to play a major role in upwelling dynamics (Veitch et al., 2010). Thus, the future of EBUS does not depend on changes of the Ekman forcing only, but also on changes of the regional ocean circulation, which is geostrophic at first order.
@@
This work focuses on the Atlantic upwelling regions which can be seen in Figure 1. In the seasonal mean from European Space Agencies Climate Change Initiative (CCI) satellite SST observations, the colder SST upwelling signal along the coast reveals regional differences. The Canary region is located from 34°N to 12°N along the Northwest-African coast. The Northern Canary Upwelling Cell (NCUC) ranging from 28°N  to 34°N is characterized by permanent weak upwelling with highest intensities in boreal summer and beginning of autumn. The Central Canary Upwelling Cell (CCUC) from 19°N to 25°N is a region of permanent strong upwelling with the highest upwelling rates in boreal summer. Figure 1 shows a larger contrast between coastal and offshore temperatures in the CCUC than in the NCUC, representing the different intensity of upwelling in the two subregions. The NCUC and CCUC summer upwelling is linked to the expansion of the Azores High and resulting wind strengthening (Arístegui et al., 2009) and is hence more pronounced in JJA (Fig. 1b). In the Southern Canary Upwelling Cell (SCUC) from 12°N to 19°N upwelling is seasonal and occurs only in boreal winter and spring (Fig. 1a), when the Inter Tropical Convergence Zone (ITCZ) moves southward towards the equator (Arístegui et al., 2009; Sylla et al., 2019). In boreal summer the ITCZ moves northward and the resulting reversal in winds favors downwelling conditions. In the summer mean the cold SST signal is therefore not visible (Fig.1b). The Benguela upwelling system is found along the Namibian and South African coast. The Northern Benguela Upwelling Cell (NBUC) experiences permanent upwelling with seasonal maxima in austral spring and winter (Fig. 1d) and spreads over the region from 16°S to 22°S (Varela et al., 2015). The strongest upwelling region in the Benguela upwelling is the Central Benguela Upwelling Cell (CBUC), including the strong Lüderitz cell, with highest upwelling rates in austral summer (Hutchings et al., 2009). In Figure 1c and 1d this can be seen by a year-round cold SST signal with high zonal extent. This region is defined from 22°S to 28°S. In the Southern Benguela Upwelling Cell (SBUC) from 28°S to 33°S upwelling occurs from austral autumn until spring, with a maximum in austral summer. While Figures 1c and 1d both show a cold SST at the coast, the contrast between coastal and offshore temperature is higher in DJF, meaning stronger upwelling rates in this season. This separation in subregions is not only important for the distinction of the seasonal cycles and different impacting mechanisms, but different regions might also be impacted differently by the global climate change.
@@
In the Canary upwelling cell, Sylla et al. (2019) used CMIP5 models to assess future changes in the Senegalo-Mauritanian upwelling region (12-20°N). They found a decrease of upwelling revealed by both SST and Ekman based indices in the southern region and unsignificant changes (Ekman index) or a decrease of upwelling (SST index) in the northern region. Recently, Jing et al. (2023) and Chang et al. (2023) evaluated future trends in upwelling regions using coupled global simulations performed with a high resolution coupled climate model (1/10° ocean resolution). Although these two studies use the same simulations, their conclusions for future trends in the Benguela region differ. While Chang et al. (2023) find negative trends for vertical velocities and positive changes in Ekman transport Jing et al. (2023) find an increase in all indices.
@@
These results highlight the necessity to consider more models for studying the oceans response to the external forcing. We will also show that partitioning the upwellings into subregions helps interpret the apparent contradictions between Jing et al. (2023) and Chang et al. (2023). Our analysis is based on vertical transport because it is a direct measure of upwelling, including geostrophic and ageostrophic components, and we introduce a new representation of the integrated transport that provides information about the characteristics of upwelling cells: strength, depth and distance to the coast.
@@
In section 2 we explain the methods and present the model and observation data used in this work. In section 3 we compute the seasonal cycle of upwelling indices during the historical period and compare them to observations in order to validate the models. The structure of the upwelling cells is analyzed in relation with the wind and SST indices, and future trends within the highest emission scenario are presented and discussed. The conclusion is drawn in section 4.
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
(d) Refer to caption
Figure 1: Seasonal mean Sea Surface Temperature from CCI Satellite data averaged over the historical 1995-2014 period for the Canary Upwelling region in boreal winter (a) and summer (b), and the Benguela Upwelling region in boreal winter (c) and summer (d). Black boxes show the subregions of each upwelling used for calculating indices.
@@ 2.1 CMIP6 model data
Table 1: CMIP6 models used in this work and their resolution (Gutiérrez et al., 2021).
HadGEM3-GC31-LL 140km, 85L 70km, 75L ([dataset] Ridley et al., 2018)
HadGEM3-GC31-MM 60km, 85L 20km, 75L ([dataset] Ridley et al., 2019)
HadGEM3-GC31-HH 30km, 85L 7km, 75L ([dataset] Roberts, 2018)
CNRM-CM6-1 140km, 91L 70km, 75L ([dataset] Voldoire, 2019b)
CNRM-CM6-1-HR 50km, 91L 20km, 75L ([dataset] Voldoire, 2019a)
@@ 2.3.2 Vertical transport and vertical transport index
Figure 2: Annual mean vertical transport $VT$ (color-filled black contours in m2 s-1) averaged in time over the historical 1995-2014 period and in latitude over the entire region retained for the: (a) Canary Upwelling; (b) Benguela Upwelling. Contours of potential temperature (solid grey in C) and mixed layer depth (solid white) are superimposed. The model shown here is HadGEM3-GC3.1-HH.
@@
$VT$ is subsequently used to construct one other single value upwelling index that may be compared to $SI$, $EI$, but that better represent the volume of waters being uplifted towards the surface within a coastal band. For consistency with $EI$, this coastal band is considered to extend as far as $1^{\circ}$ offshore $X_{1^{\circ}}$ (Tim et al., 2015). The Vertical Transport Index ($VTI$) is additionally defined as the vertical transport $VT$ at a fixed depth of 50 m (Jing et al., 2023):
@@ 3.1 Influence of model resolution on the SST bias
For all models except HadGEM3-GC3.1-HH, SST biases are found to be much higher in the Benguela region than in the Canary. Within the Canary system, biases do not exceed +2C in the subtropics (lat>20N), and may even become sightly negative in the tropics. Within the Benguela system, biases exceed 3C in the subtropics, and gradually increase towards the equator until reaching 9C in the tropical southern Atlantic. Although one would expect models with a higher resolution to perform better in representing the SST (Park and Latif, 2020), increasing the model’s resolution did not systematically improve the model’s SST in the 5 simulations considered here. Within the Canary system, as resolution gets higher, the warm subtropical bias is reduced in the subtropics, but the cold bias can get worse in the tropics. Within the Benguela, the warm bias remains as high as in the 20 km resolution model as in the 70 km models at all latitudes. Nevertheless, the HadGEM3-GC3.1-HH shows an above average performance in regards to the southern Benguela region, with a maximum positive bias of 3°C in the northern region and a slight underestimation of SST of up to 1°C in the southern region. This may be seen as a remarkable performance as such bias is lower than the one of the 1/10°resolution Community Earth System Model High resolution model (CESM-H) (Jing et al., 2023; Chang et al., 2023). Richter and Tokinaga (2020) already found that the HadGEM family to be one of the best performing in the reduction of the Tropical Atlantic SST error.
@@
Figure 3: Annual mean SST bias showed by the CMIP6 models over the 1995-2014 period in the Atlantic EBUS: (a) Canary; (b) Benguela. $SST_{\mbox{bias}}=SST_{\mbox{CMIP6}}-SST_{\mbox{CCI}}$.
@@ 3.2 Seasonal cycle
In the past, it has been found that according to the choice of the season upwelling intensity can differ (Varela et al., 2015) and that models do not always capture upwelling onset correctly (Richter and Tokinaga, 2020; Sylla et al., 2019). In the Atlantic, the southernmost regions of the Canary and Benguela upwellings have specific seasonal dynamics (Strub et al., 2013; Sylla et al., 2019). We therefore assess the climatologies of the seasonal cycle using the subregions defined in Figure 1. We compare the models to observation data for the surface upwelling indices which helps to further evaluate the models performance. We then introduce the $VTI$ derived from vertical water mass transport and compare it to the surface indices. The seasonal cycle for the three indices and for each of the subregions can be seen in Figure 4 for the Canary Upwelling System and in Figure 5 for the Benguela Upwelling System. The seasonal cycle of the SST index from the models and CCI satellite observation is shown in panels a,d,g. In the NCUC HadGEM3-GC3.1-LL shows a negative SST index in boreal summer (Figure 4a). This is due to strong coastal warming along the Moroccan coast in this model during the summer months. The reasons for this have not been explored further. Apart from this model, the models display the onset of upwelling in early summer in the NCUC compared to CCI, although underestimating its intensity. In the CCUC (Fig. 4d) all models portray the year round positive SST Index also found in CCI, but fail to represent its intensity and its maxima. However, in Figure 4g we see that upwelling onset is portayed rather well, where in previous work using the SST index from CMIP5 models in the SCUC, the models (notably the previous generation of the two models used here) have failed to reproduce the seasonal cycle and the onset of upwelling (Sylla et al., 2019). In the Benguela Upwelling System (Fig. 5) the SST Index (panels a,d,g) shows negative values for several models, especially in the NBUC (5a). This is linked to the Tropical Atlantic SST bias (see section 3.1, Fig. 3), which is more intense at the coast. As expected from the analysis of the SST bias, higher resolution models also perform better in depicting the SST index than the lower resolution members of the same family, with HadGEM3-GC3.1-HH being the closest to observations. Upwelling onset is delayed compared to CCI, with the delay becoming more pronounced as the the index becomes lower.
@@
While for the SST index the models show a large difference in terms of amplitude and upwelling onset, the Ekman index (panels b,e,h in Fig. 4 and 5) is closer to the observations in both upwelling systems. This has already been found in previous work in the SCUC (Sylla et al., 2019). The Ekman Index follows the same seasonal cycle and the differences between the models’ performance are small. Nonetheless, the Ekman transport is overestimated in the NBUC in austral winter by most models compared to observations. Previous work has shown that an improvement of model SST in this region often goes along with overestimation of wind (Richter and Tokinaga, 2020). This is consistent with the overestimation in HadGEM3-GC3.1-HH winds in relation with its comparably good performance in the SST index in the Benguela upwelling system. However, this does not explain in the overestimation in the two CNRM models. HadGEM3-GC3.1-MM shows a very good performance in all three subregions in the Benguela for the $EI$. When comparing the Ekman index with the SST index an obvious difference between the seasonal cycle and the onset of upwelling becomes evident. This is the case for modeled as much as the observed data. Upwelling onset is observed later in the season for the SST index than for the other indices. A later onset of upwelling from a SST derived compared to an Ekman divergence derived index has also been observed by Sylla et al. (2019).
Figure 4: Mean monthly upwelling indices over the time period 1995-2014 and their standard deviation for the three subregions in the Canary Upwelling region for the five CMIP6 models and the SST observations from CCI (panels a,d,g) and the ERA5 Reanalysis wind stress (panels b,e,h).
@@
When comparing SST and Ekman indices to the Vertical Transport index (panels c,f,i), the seasonal cycle from $VTI$ appears very close in intensity and phase to the $EI$. The $EI$ can be compared directly to $VTI$, as they have the same units. The seasonal cycle of the $VTI$ behaves very similar to the $EI$ for both upwelling regions and all subregions. This shows that there is a direct oceanic response to the wind forcing, which is reproduced by the models dynamics. Apart from a few extrema in the Benguela Upwelling Region (Fig. 5), the CNRM models tend to show lower vertical transport rates than the HadGEM models. This is in accordance with the $EI$. The difference in intensity between $EI$ and $VTI$ is highest the closest to the equator in both regions (Fig. 4i and Fig. 5c), while in the poleward portions (Fig. 4c and Fig. 5i) intensities and seasonal cycle of $EI$ and $VTI$ correspond almost perfectly. This indicates that other transport processes might be more important in the equatorward portions of the upwellings. This could be linked to the convergence of surface flows within the coastal band, onshore or alongshore, due to the large-scale circulation (Veitch et al., 2010).
Figure 5: Mean monthly upwelling indices over the time period 1995-2014 and their standard deviation for the three subregions in the Benguela Upwelling region for the five CMIP6 models and the SST observations from CCI (panels a,d,g) and the ERA5 Reanalysis wind stress (panels b,e,h).

### v2

@@ Abstract
Eastern Boundary Upwelling Systems (EBUS) are characterized by wind-triggered upwelling of deep waters along the coast. They are hotspots of biological productivity and diversity and therefore have a high economic, ecological and social importance. In the past, different methods using surface data have been used to estimate upwelling. Recently, the IPCC has suggested directly assessing vertical velocities as a promising method. We use this method to study the two Atlantic EBUS from CMIP6 models from the HadGEM3-GC3.1 and the CNRM6-CM6 family, for both the historical period and a high-emission future scenario with spatial resolutions in the ocean component ranging from 1°to 1/12°. The two major upwelling regions are divided in subregions depending on their seasonality. The vertical transport index shows similar values to a wind-derived Ekman index. Directly evaluating upwelling from transport processes further provides information about the depth of the upwelling, which has previously been identified as an important factor for nutrient availability. We show that depending on the subregion of the upwelling system, different cell structures can be seen in terms of depth and distance to the coast of maximum velocities. When looking at possible future changes, high interannual variability limits the significance of the trends but could indicate a poleward shift of the upwelling regions. A detailed comparison of the spatial structures and the distinction in subregions is important to explain contradictory trends in previous works.
@@ 1 Introduction
Within the context of climate change, considering the future of EBUS therefore becomes of primary importance. Given that the major drivers of the upwelling in EBUS are the along-shore trade winds, the community has focused its efforts in considering projected wind changes. Two main hypotheses are usually considered: i) the Bakun hypothesis (Bakun, 1990) states that along-shore wind stress will increase in future climates because of the intensification of the inland low pressure systems on the eastern borders of these regions. Seawater heat capacity being higher than soil heat capacity, one expects the surface inland temperature to rise faster than the seawater one, and therefore the atmospheric thermal low over land to strengthen. ii) The second hypothesis states that there will be a poleward migration of the trade winds as the atmospheric Hadley cell expands poleward, driving in turn the poleward shift of the subtropical high pressure atmospheric cells found in the four ocean basins (Rykaczewski et al., 2015). While one expects a general increase of the volume of water that is being upwelled to the surface in the case of an overall increase of the trade winds (hypothesis i), along-shore regional disparities are expected to happen under the poleward migration hypothesis, with an intensification/decrease of the upwelling in the poleward/equatorward parts of the EBUS.
Gulev et al. (2021)’s recent analysis, based on historical datasets of the wind fields, led to a rejection of the Bakun hypothesis. Although the increase of surface temperature over the last century was higher over land (1.59 [1.34 to 1.83]°C) than over the ocean (0.88 [0.68 to 1.01]°C), this increase did not significantly impact the intensity of the upwelling favourable surface wind stress. On the other hand, the poleward migration of the Hadley cell was confirmed, and found to be more pronounced in the Northern than in the Southern Hemisphere. Additionally, contradicting trends were found for different EBUS. However, as already shown in the work of Sydeman et al. (2014) and Varela et al. (2015), trends appear to be very sensitive to the data product, period and geographical boundaries of the region of interest. This sensitivity highlights the lack of long-term robust observational datasets that can be used by the community to infer trends in upwelling systems.
Further, the ocean circulation, forced in part by the wind stress curl, is known to play a major role in upwelling dynamics (Veitch et al., 2010). Thus, the future of EBUS does not depend on changes of the Ekman forcing only, but also on changes of the regional ocean circulation, which is geostrophic to first order.
@@
This work focuses on the Atlantic upwelling regions which can be seen in Fig. 1. In the seasonal mean from European Space Agencies Climate Change Initiative (CCI) satellite SST observations, the colder SST upwelling signal along the coast reveals regional differences. The Canary region is located from 34°N to 12°N along the Northwest-African coast. The Northern Canary Upwelling Cell (NCUC) ranging from 28°N  to 34°N is characterized by permanent weak upwelling with highest intensities in boreal summer and beginning of autumn. The Central Canary Upwelling Cell (CCUC), from 19°N to 25°N  is a region of permanent strong upwelling with the highest upwelling rates in boreal summer. Figure 1 shows a larger contrast between coastal and offshore temperatures in the CCUC than in the NCUC, representing the different intensity of upwelling in the two subregions. The NCUC and CCUC summer upwelling is linked to the expansion of the Azores High and resulting wind strengthening (Arístegui et al., 2009) and is hence more pronounced in JJA (Fig. 1b). In the Southern Canary Upwelling Cell (SCUC), from 12°N to 19°N, upwelling is seasonal and occurs only in boreal winter and spring (Fig. 1a), when the Inter Tropical Convergence Zone (ITCZ) moves southward towards the equator (Arístegui et al., 2009; Sylla et al., 2019). In boreal summer, the ITCZ moves northward and the resulting reversal in winds favors downwelling conditions. The cold SST signal is therefore not visible in the summer mean (Fig.1b). The Benguela upwelling system is found along the Namibian and South African coast. The Northern Benguela Upwelling Cell (NBUC) experiences permanent upwelling with seasonal maxima in austral spring and winter (Fig. 1d) and spreads over the region from 16°S to 22°S (Varela et al., 2015). The strongest upwelling region in the Benguela upwelling is the Central Benguela Upwelling Cell (CBUC). It includes the strong Lüderitz cell, and has the highest upwelling rates in austral summer (Hutchings et al., 2009). In Fig. 1, Panels c and d, this can be seen by a year-round cold SST signal with high zonal extent. This region is defined from 22°S to 28°S. In the Southern Benguela Upwelling Cell (SBUC), from 28°S to 33°S, upwelling occurs from austral autumn until spring, with a maximum in austral summer. While Fig. 1, Panels c and d, both show a cold SST at the coast, the contrast between coastal and offshore temperature is higher in DJF, meaning stronger upwelling rates in this season. This separation in subregions is not only important for the distinction of the seasonal cycles and different impacting mechanisms, but different regions might also be impacted differently by the global climate change.
@@
In the Canary upwelling cell, Sylla et al. (2019) used CMIP5 models to assess future changes in the Senegalo-Mauritanian upwelling region (12-20°N). They found a decrease of upwelling revealed by both SST and Ekman based indices in the southern region and unsignificant changes (Ekman index) or a decrease of upwelling (SST index) in the northern region. Recently, Jing et al. (2023) and Chang et al. (2023) evaluated future trends in upwelling regions using coupled global simulations performed with a high resolution coupled climate model (1/10° ocean resolution). Although these two studies use the same simulations, their conclusions for future trends in the Benguela region differ. While Chang et al. (2023) find negative trends for vertical velocities and positive changes in Ekman transport, Jing et al. (2023) find an increase in all indices.
@@
These results highlight the necessity to consider more models for studying the oceans response to the external forcing. We will also show that partitioning the upwellings into subregions helps interpret the apparent contradictions between Jing et al. (2023) and Chang et al. (2023). Our analysis is based on vertical transport because it is a direct measure of total upwelling. It includes the contribution of all physical processes creating upwelling. We introduce a new representation of the integrated transport that provides information about the characteristics of upwelling cells: strength, depth and distance to the coast.
@@
In section 2, we explain the methods and present the model and observational data used in this work. In section 3, we compute the seasonal cycle of upwelling indices during the historical period and compare them to observations in order to validate the models. The structure of the upwelling cells is analyzed in relation with the wind and SST indices, and future trends within the highest emission scenario are presented and discussed. The conclusion is drawn in section 4.
@@ 2.3.2 Vertical transport and vertical transport index
$VT$ is subsequently used to construct one other single value upwelling index that may be compared to $SI$ and $EI$, but that better represents the volume of waters being uplifted towards the surface within a coastal band. For consistency with $EI$, this coastal band is considered to extend as far as $1^{\circ}$ offshore $X_{1^{\circ}}$ (Tim et al., 2015). The Vertical Transport Index ($VTI$) is additionally defined as the vertical transport $VT$ at a fixed depth of 50 m (Jing et al., 2023):
@@ 3.1 Influence of model resolution on the SST bias
For all models except HadGEM3-GC3.1-HH, SST biases are found to be much higher in the Benguela region than in the Canary. Within the Canary system, biases do not exceed +2C in the subtropics (lat>20N), and may even become sightly negative in the tropics. Within the Benguela system, biases exceed 3C in the subtropics, and gradually increase towards the equator until reaching 9C in the tropical southern Atlantic. Although one would expect models with a higher resolution to perform better in representing the SST (Park and Latif, 2020), increasing the model’s resolution did not systematically improve the model’s SST in the 5 simulations considered here. Within the Canary system, as resolution gets higher, the warm subtropical bias is reduced in the subtropics, but the cold bias can get worse in the tropics. Within the Benguela, the warm bias remains as high in the 20 km resolution model as in the 70 km models at all latitudes. Nevertheless, the HadGEM3-GC3.1-HH shows an above average performance in regards to the southern Benguela region, with a maximum positive bias of 3°C in the northern region and a slight underestimation of SST of up to 1°C in the southern region. This may be seen as a remarkable performance as such bias is lower than the one of the 1/10°resolution Community Earth System Model High resolution model (CESM-H) (Jing et al., 2023; Chang et al., 2023). Richter and Tokinaga (2020) already found that the HadGEM family to be one of the best performing in the reduction of the Tropical Atlantic SST error.
@@ 3.2 Seasonal cycle
In the past, it has been found that models do not always capture the seasonal upwelling onset correctly (Varela et al., 2015; Richter and Tokinaga, 2020; Sylla et al., 2019). In the Atlantic, the southernmost regions of the Canary and Benguela upwellings have specific seasonal dynamics (Strub et al., 2013; Sylla et al., 2019). We therefore assess the climatologies of the seasonal cycle using the subregions defined in Fig. 1. We compare the models to observational data for the surface upwelling indices which helps to further evaluate the model’s performance. We then introduce the $VTI$ derived from vertical water mass transport and compare it to the surface indices. The seasonal cycle for the three indices and for each of the subregions can be seen in Fig. 4 for the Canary Upwelling system and in Fig. 5 for the Benguela Upwelling system. The seasonal cycle of the SST index from the models and CCI satellite observation are shown in Fig. 4, Panels a,d,g. In the NCUC HadGEM3-GC3.1-LL shows a negative SST index in boreal summer (Fig. 4a). This is due to strong coastal warming along the Moroccan coast in this model during the summer months. The reasons for this have not been explored further. Apart from this model, the models display the onset of upwelling in early summer in the NCUC compared to CCI, although underestimating its intensity. In the CCUC (Fig. 4d) all models portray the year round positive SST Index also found in CCI, but fail to represent its intensity and its maxima. However, in Fig. 4g we see that upwelling onset is portayed rather well, where in previous work using the SST index from CMIP5 models in the SCUC, the models (notably the previous generation of the two models used here) have failed to reproduce the seasonal cycle and the onset of upwelling (Sylla et al., 2019). In the Benguela Upwelling System, the SST Index shows negative values for several models, especially in the NBUC (Fig. 5, Panels a,d,g). This is linked to the Tropical Atlantic SST bias (see section 3.1, Fig. 3), which is more intense at the coast. As expected from the analysis of the SST bias, higher resolution models also perform better in depicting the SST index than the lower resolution members of the same family, with HadGEM3-GC3.1-HH being the closest to observations. Upwelling onset is delayed compared to CCI, with the delay becoming more pronounced as the the index becomes lower.
@@
While for the SST index the models show a large difference in terms of amplitude and upwelling onset, the Ekman index (Figs. 4 and 5, Panels b, e, h) is closer to the observations in both upwelling systems. This has already been found in previous work in the SCUC (Sylla et al., 2019). The Ekman Index follows the same seasonal cycle and the differences between the models’ performance are small. Nonetheless, the Ekman transport is overestimated in the NBUC in austral winter by most models compared to observations. Previous work has shown that an improvement of model SST in this region often goes along with overestimation of wind (Richter and Tokinaga, 2020). This is consistent with the overestimation in HadGEM3-GC3.1-HH winds in relation with its comparably good performance in the SST index in the Benguela upwelling system. However, this does not explain the overestimation in the two CNRM models. HadGEM3-GC3.1-MM shows a very good performance in all three subregions in the Benguela for the $EI$. When comparing the Ekman index with the SST index an obvious difference between the seasonal cycle and the onset of upwelling becomes evident. This is the case for modeled as much as the observed data. Upwelling onset is observed later in the season for the SST index than for the other indices. A later onset of upwelling from a SST derived compared to an Ekman divergence derived index has also been observed by Sylla et al. (2019).
@@
When comparing SST and Ekman indices to the Vertical Transport index (Panels c,f,i), the seasonal cycle from $VTI$ appears very close in intensity and phase to the $EI$. The $EI$ can be compared directly to $VTI$, as they have the same units. The seasonal cycle of the $VTI$ behaves very similar to the $EI$ for both upwelling regions and all subregions. This shows that there is a direct oceanic response to the wind forcing, which is reproduced by the models dynamics. Apart from a few extrema in the Benguela Upwelling Region (Fig. 5), the CNRM models tend to show lower vertical transport rates than the HadGEM models. This is in accordance with the $EI$. The difference in intensity between $EI$ and $VTI$ is highest the closer to the equator in both regions (Fig. 4i and Fig. 5c), while in the poleward portions (Fig. 4c and Fig. 5i) intensities and seasonal cycle of $EI$ and $VTI$ correspond almost perfectly. This indicates that other transport processes might be more important in the equatorward portions of the upwellings. This could be linked to the convergence of surface flows within the coastal band, onshore or along-shore, due to the large-scale circulation (Veitch et al., 2010).